OpenStudy (anonymous):

If x+y+z=0 and x^2+y^2+z^2=7, then x^2y^2+y^2z^2+z^2x^2=a/b, where a and b are positive coprime integers. What is a+b?

OpenStudy (anonymous):

Just kidding.. :)

OpenStudy (anonymous):

What ? o.O

OpenStudy (anonymous):

we have \[ (x+y+z)^2=0\\ \sum x+2\sum xy=0\\ xy+yz+zx=0 \]

OpenStudy (anonymous):

similarly, \[ (xy+yz+zx)^2=0\\ \sum(xy)^2+2\sum(xy)(yz)=0\\ {a\over b}=-2(xy^2z+zyz^2+x^yz)=-2(xyz)(x^2+y^2+z^2) \]

terenzreignz (terenzreignz):

Wondering how you could get \(\large xy + yz + zx = 0\) \[\large (x+y+z)^2 = 0 = x^2+y^2+z^2+2xy+2yz+2zx\] \[\large 0 = 7 + 2xy + 2yz+2zx\] \[\huge -\frac72= xy+yz+zx\]

OpenStudy (anonymous):

doh! :P

OpenStudy (anonymous):

but thats the logic.

terenzreignz (terenzreignz):

This question is not good for my health... I'm getting nauseous :D

OpenStudy (anonymous):

lol i got that -7/2 result too :p

terenzreignz (terenzreignz):

Time to toss ideas around... \[\large (2xy+2yz+2zx)^2=4(x^2y^2+y^2z^2+z^2x^2)+8x^2yz+8xy^2z+8xyz^2\]

terenzreignz (terenzreignz):

bloody hell \[\ (2xy+2yz+2zx)^2=4(x^2y^2+y^2z^2+z^2x^2)+8x^2yz+8xy^2z+8xyz^2\]

OpenStudy (anonymous):

\[ x+y+z=0 \] \[ x^{2}+y^{2}+z^{2}=7 \] \[ x^{2}y^{2}+y^{2}z^{2}+z^{2}x^{2}=\frac{a}{b}\mbox{, a and b positive and coprime} \] find \(a+b\) \[ (x+y+z)^{2}=x^{2}+y^{2}+z^{2}+2xy+2yz+2zx=0 \] \[ xy+yz+zx=-\frac{7}{2} \] Squaring both sides \[ x^{2}y^{2}+y^{2}z^{2}+z^{2}x^{2}+2x^{2}yz+2xy^{2}z+2xyz^{2}=\frac{49}{4} \] \[ x^{2}y^{2}+y^{2}z^{2}+z^{2}x^{2}+2(xyz)(x+y+z)=\frac{49}{4} \] since \(x+y+z=0\) \[ x^{2}y^{2}+y^{2}z^{2}+z^{2}x^{2}=\frac{49}{4} \]

OpenStudy (anonymous):

49/4 can't be simplified any further and 49 and 4 are coprime so a + b = 53

terenzreignz (terenzreignz):

I cannot believe I missed that , @Meepi :)

OpenStudy (anonymous):

hey .. how about a medal to get you guys started???

OpenStudy (anonymous):

:D

terenzreignz (terenzreignz):

Now, where is @vikrantg4 ?

OpenStudy (anonymous):

I was thinking the same but I thought that I would not make it to the final answer...

OpenStudy (anonymous):

LoL meepi rocked !